2 g of a substance dissoled in 40 g of water produced a depression of 274.5 k in the freezing point of water. Calculate the molecular mass of the substance. The molal depression constant for water is 274.85 k per molal.

Grade:12

1 Answers

Labanya
6 years ago
w2(solute)=2g w1(water)= 40g kf= 274.85 kkgmol^-1 delTf= 274.5k m2=?therefore, we knowdelTf=kf×molalitydelTf= kf× 1000×w2/(M2*w1)M2=kf*1000*w2/(delTf*w1)M2= 274.85×1000×2/(274.5×40)M2= 50.06 gmol^-1